How many grams of Sodium Hydroxide is needed to produce 242g of table salt (NaCl)?

Fe + NaOH ---> Fe (OH)3 +NaCl


1
Expert's answer
2021-02-02T01:22:02-0500

FeCl3 + 3NaOH ---> Fe (OH) 3 + 3NaCl

"\nu" (NaCl) = m/M= 242/58,5=4.14 mol

"\nu" (NaCl)/3="\nu"(NaOH)/3"\implies" "\nu"(NaOH)="\nu" (NaCl)= 4.14 mol

m(NaOH)="\nu"*M=4.14*40=165.6 g